# Experimental review on the chiral magnetic effect in relativistic heavy ion collisions

### v2

@@ Abstract
The chiral magnetic effect (CME) refers to a predicted phenomena in quantum chromodynamics that manifests as a charge separation along an external magnetic field, driven by an imbalance of quark chirality. Searches for the CME has been carried out by azimuthal particle correlations in relativistic heavy ion collisions where such a chirality imbalance is anticipated and a strong magnetic field is created in the initial stage. No conclusive experimental evidence on the CME has been established so far because of large background contributions to azimuthal correlation observables. We review the status of the experimental search for the CME, covering the observables used, the techniques to mitigate backgrounds, and the strengths and limitations of various experimental approaches, and outline a future prospect of the CME search in high-energy nuclear collisions.
@@ 1 Introduction
Relativistic heavy ion collisions provide an ideal platform to search for the CME. The high temperatures achieved in these collisions can significantly enhance the vacuum transition probabilities [6, 7, 3]. A strong magnetic field on the order of $B\sim m_{\pi}^{2}/e\sim 10^{18}$ Gauss is believed to be created in non-central heavy ion collisions (here $e$ is the elementary charge and $m_{\pi}$ is the pion mass) [8, 9, 10]. This magnetic field is expected to quickly fade off as the two spectator remnants recede from each other. However, the decay time can be significantly prolonged depending on the electric conductivity of the QGP [11, 12, 13, 14, 15]. With plausible parameters for the axial current density and the time evolution of the magnetic field, anomalous viscous fluid dynamics (avfd) calculations suggest that an observable CME signal can indeed manifest itself in relativistic heavy ion collisions [16, 17, 18]. On the other hand, a significantly smaller signal than experimentally perceptible has also been theoretically suggested [19].
Experimental searches for the CME in relativistic heavy ion collisions at RHIC and the LHC have persisted over the last two decades. Significant progress has been made, while many challenges related to the treatment of background effects remain open. Here, we review the experimental status of the CME search and discuss prospects of a potential CME discovery in forthcoming high-statistics heavy ion data sets. For further reading, the reader is referred to extensive reviews on the subject in Refs. [20, 21, 22, 23, 24, 25, 26, 27, 28, 29, 30, 31, 32].
@@ 2.1 The $\Delta\gamma$ correlator
The RP and the impact parameter are theoretical concepts and can only be rigorously defined under the limit of smooth nuclear density distributions. Because nucleon distributions in nuclei are lumpy, the symmetry planes relevant for collective flow are determined by the geometry of the participating nucleons [37]. They are also called participant planes (PP) or flow planes, and they fluctuate event by event, differing for each harmonic order. Because the magnetic field is determined mostly by spectator protons, its direction is expected to be roughly perpendicular to the first-order symmetry plane, which is often called the spectator plane (SP). Thus, the $\psi_{\textsc{RP}}$ in the various terms of Eq. (1) should be replaced by the corresponding-order harmonic planes ($\psi_{n}$). In particular, that for the CME sine term can be replaced by the first-order symmetry plane, which is most closely correlated to the RP. It can also be replaced by the second-order participant plane ($\psi_{2}^{\textsc{pp}}$) because it is closely related to the RP and the most accurately assessed in experiment; in this case, the coefficients $a_{\pm}$ would be reduced accordingly.
@@ 2.2 Charge-dependent physics backgrounds
Many physics processes are intrinsically charge dependent; examples are resonance decays and (mini-)jet correlations. There are more contributions to OS pairs than to SS pairs from resonance decays. Charge-ordering is well known in jet fragmentation that also leads to an excess of OS over SS correlations. These short-range charge-dependent correlations can be generally characterized as “clusters”. Local charge conservation (LCC) is preserved in cluster decays, so these background contributions are often referred to as LCC effects.
@@ 2.3 Experimental and analysis methods
The experimentally measured CME observables may contain contributions from both the signal and various background sources. In general, one can write
@@
Because the background and signal components are entangled, a central strategy of CME analyses is to vary one component while keeping the other approximately fixed, and then study the corresponding change in the observable.
@@ 2.3.2 Event-shape engineering
On the other hand, the statistical fluctuations in the apparent event-by-event anisotropy quantities, e.g. the $v_{2}^{\text{obs}}\equiv\langle{\cos(2\phi-2\psi_{\textsc{EP}})}\rangle$ [72]) and the $q_{2}$ of single particles as well as particle pairs [73, 74], are large. It is tempting to use these quantities calculated from the POIs themselves to select events where the background contribution is small due not only to the small(er) ellipticity of the initial geometry but also to statistical fluctuations in particle distribution in a given event. The so-called event-shape selection (ESS) method is based on this idea. In such selected events the mean $v_{2}$ of POIs is dominated by statistical fluctuations because the event-selection variable is based on POIs. In events where the overall POIs’ $v_{2}=0$, the $v_{2}$ values of the background-contributing sources are unnecessarily zero [75]. Such biases are difficult to disentangle, as it would inevitably require the full knowledge of all underlying components of the event. Further quantitative investigations are required to clarify the practical applicability of the ESS method in CME searches.
@@ 2.3.4 Comparative measurements with spectator/participant planes
The elliptic flow values measured with respect to the PP, $v_{2}^{\textsc{pp}}$, and to the SP, $v_{2}^{\textsc{sp}}$, differ by an appeciable 10-20% depending on the collision centrality. This provides an opportunity to extract the CME signal fraction from the two $\Delta\gamma$ measurements with respect to SP and PP, in place of $\psi_{\textsc{rp}}$ in Eq. (2), by [83, 77]
@@ 3.2 Evidence of backgrounds
The first unambiguous experimental evidence of backgrounds dominating the $\Delta\gamma$ observable is from measurements in small systems, where the CME signal is expected to be negligible. Such measurements have been performed in p+Pb collisions by the CMS Collaboration at the LHC [81], and in pp+Au and dd+Au collisions by the STAR Collaboration at RHIC [99]. These results are shown in Fig. 5, where large differences between $\gamma_{\textsc{os}}$} and $\gamma_{\textsc{ss}}$} are observed, comparable to those measured in peripheral heavy ion collisions at similar multiplicities. Because the magnetic field direction is not correlated with the participant plane direction in small systems, which arises from pure fluctuation effects, any CME signals, even existent in those small-system collisions, are not observable by the $\Delta\gamma$ correlator [81, 82]. The observed $\Delta\gamma$ in small systems and peripheral heavy ion measurements are likely dominated by RP-independent three-particle correlations, as suggested by hijing model calculations [100].
@@ 3.3.1 Event-shape engineering
It is worthwhile to note that all ESE analyses assumed that the effects of nonflow and RP-independent three-particle correlation backgrounds are negligible. These effects will have to be accounted for in future high-statistics measurements.
One shortcoming of the ESE method is the extrapolation to zero $v_{2}$, which causes large uncertainties, because of the limited range of dynamical fluctuations of $v_{2}$. It is attempting to exploit statistical fluctuations of $v_{2}$ by using the same POIs as a means to select events. Such an attempt was first reported by the STAR Collaboration in Ref. [72]. Figure 9 presents an example of such an analysis. Unfortunately, such an analysis does not fully eliminate the flow-induced background because the average $v_{2}$’s of background contributing sources are no longer proportional to the statistically fluctuating $v_{2}$ of the POIs used in the analysis [75]. The latest attempt in this ESS approach is to use the $q_{2}$ of pairs of POIs for the event-shape selection. Example results from such an attempt is shown in Fig. 9 [74, 69]. The pair $q_{2}$ and $v_{2}$, although directly containing those from the background contributing sources, are dominated by single-particle level contributions from POIs and contains a self-correlation between pairs sharing a common POI [50, 107]. Because the POI pair $q_{2}$ is still selecting on the statistical fluctuations of $v_{2}$, it is challenging, if not at all impossible, to determine how much background remains in the intercept [73, 107]. Further quantitative studies and modeling will be required.
@@ 3.3.2 Isobar collisions
The Ru+Ru to Zr+Zr ratio of the $\Delta\gamma/v_{2}$ variable was observed to be smaller than unity (see Fig. 10), contrary to the expectation of a larger CME signal in Ru+Ru and the same background contributions. The smaller-than-unity ratio results from a few percent difference in the multiplicities, greater (and thus smaller background contribution, see Eq. (9)] in Ru+Ru than Zr+Zr collisions. Such a difference in multiplicities was predicted by energy density functional theory calculations to root in the difference in nuclear structure between the $^{96}_{44}$Ru and $^{96}_{40}$Zr nuclei [84, 110, 111, 112]: the larger number of neutrons in $^{96}_{40}$Zr yields a thicker neutron skin, larger overall size, smaller energy density, and hence lower multiplicity in Zr+Zr collisions.
@@
Such an estimate of the “baseline” in the isobar comparison has been made by the STAR Collaboration in Refs. [113, 114]. A 2D fitting method of two-particle correlations in $(\Delta\eta,\Delta\phi)$ was used to estimate the (difference) in nonflow contamination in $v_{2}{\{2\}}$ measurements. The three-particle nonflow contamination was estimated by the hijing model following Ref. [115]. Since hijing does not include collective flow, its entire three-particle correlations can be used as an estimate of the RP-independent three-particle correlation background [117, 118]. It was found that the nonflow contribution to $v_{2}$ and the three-particle correlation background were both smaller in Ru+Ru collisions, presumably due to the larger multiplicity dilution in Ru+Ru collisions. The two effects partially cancel each other, and the net effect turns out to be slightly negative compared to the baseline estimated from $r$. These are shown by the brown lines accompanied by shaded bars in Fig. 10 for four of the $\Delta\gamma/v_{2}$ isobar ratio measurements, where the cumulant method is used in measuring $v_{2}{\{2\}}$. The measurements of the $\Delta\gamma/v_{2}$ isobar ratios are consistent with the estimated background baselines, within approximately $1\sigma$ uncertainty. An upper limit of approximately 10% for $f_{\textsc{cme}}$ is extracted for each of the four measurements at the confidence level of 95%. For reference, the highest precision data point of the $\Delta\gamma/v_{2}$ isobar ratios in Fig. 10 has a statistical uncertainty of 0.4%, which corresponds to an uncertainty of approximately $3\%$ in terms of $f_{\textsc{cme}}$.
@@ 3.3.3 Spectator/participant planes
As mentioned in Ref. [122], the $f_{\textsc{cme}}^{\text{obs}}$ measurements are still contaminated by nonflow correlations; see Sect. 2.2.2. These nonflow correlations include two-particle nonflow in the $v_{2}{\{2\}}$ measurement and RP-independent three-particle correlations in the $C_{3}$ measurement. Both affect only measurements relative to the PP inferred from TPC particles, not those relative to the SP measured by the ZDCs. The nonflow contribution to $v_{2}{\{2\}}$ reduces $\Delta\gamma\{{\textsc{tpc}}\}$ resulting an overestimate of the ratio $\Delta\gamma^{\textsc{sp}}/\Delta\gamma^{\textsc{pp}}$, thus a positive contribution to the $f_{\textsc{cme}}^{\text{obs}}$ estimate. The RP-independent three-particle contamination increases $\Delta\gamma\{{\textsc{tpc}}\}$ making its difference from $\Delta\gamma\{{\textsc{zdc}}\}$ smaller, thus gives a negative contribution to the $f_{\textsc{cme}}^{\text{obs}}$ estimate. These two effects partially cancel each other.
The nonflow and the RP-independent three-particle correlation effects in SP/PP measurements have been investigated in Ref. [115] using models. The nonflow effect in $v_{2}$ was estimated by using the ampt model. The effect of RP-independent three-particle correlations is estimated by using the hijing model. The estimated nonflow contribution to $f_{\textsc{cme}}^{\text{obs}}$ is shown in Fig. 12 by the points connected by lines, compared to the STAR measurement. It is found that the net effect of nonflow correlations causes a nearly zero or perhaps negative bias to $f_{\textsc{cme}}$. It will be interesting to have a more data-driven analysis of background effects in the future.

### v3

@@ Abstract
The chiral magnetic effect (CME) refers to a predicted phenomenon in quantum chromodynamics that manifests as a charge separation along an external magnetic field, driven by an imbalance of quark chirality. Searches for the CME have been carried out by azimuthal particle correlations in relativistic heavy ion collisions where such a chirality imbalance is anticipated and a strong magnetic field is created in the initial stage. No conclusive experimental evidence on the CME has been established so far because of large background contributions to azimuthal correlation observables. We review the status of the experimental search for the CME, covering the observables used, the techniques to mitigate backgrounds, and the strengths and limitations of various experimental approaches, and outline a future prospect of the CME search in high-energy nuclear collisions.
@@ 1 Introduction
Relativistic heavy ion collisions provide an ideal platform to search for the CME. The high temperatures achieved in these collisions can significantly enhance the vacuum transition probabilities [6, 7, 3]. A strong magnetic field on the order of $B\sim m_{\pi}^{2}/e\sim 10^{18}$ Gauss is believed to be created in non-central heavy ion collisions (here $e$ is the elementary charge and $m_{\pi}$ is the pion mass) [8, 9, 10]. This magnetic field is expected to quickly fade off as the two spectator remnants recede from each other. However, the decay time can be significantly prolonged depending on the electric conductivity of the QGP [11, 12, 13, 14, 15]. With plausible parameters for the axial current density and the time evolution of the magnetic field, anomalous viscous fluid dynamics (avfd) calculations suggest that an observable CME signal could indeed manifest itself in relativistic heavy ion collisions [16, 17, 18]. On the other hand, a significantly smaller signal than experimentally perceptible has also been theoretically suggested [19].
Experimental searches for the CME in relativistic heavy ion collisions at RHIC and the LHC have persisted over the last two decades. Significant progress has been made, while many challenges related to the treatment of background effects remain open. Here, we review the experimental status of the CME search and discuss prospects of a potential CME discovery in forthcoming high-statistics heavy ion data sets. For further reading, the reader is referred to extensive reviews on the subject in Refs. [20, 21, 22, 23, 24, 25, 26, 27, 28]. We note that this is an experimental review so the theoretical description/discussion is kept minimal, mostly at the level needed to interpret observables. For in-depth theoretical discussions, the reader is referred to extensive literature on the subject, for example, Refs. [29, 30, 31, 32, 33, 34].
The rest of the review is organized as follows. Section 2 describes the main observable used in most of the CME searches, the physics backgrounds associated with the observable, and experimental ways to mitigate these backgrounds. Section 3 presents early measurements, discusses the physics backgrounds in these measurements, and highlights the recent measurements where these backgrounds are mitigated. Finally, a summary and outlook is given in Sect. 4.
@@ 2.1 The $\Delta\gamma$ correlator
The RP and the impact parameter are theoretical concepts and can only be rigorously defined under the limit of smooth nuclear density distributions. Because nucleon distributions in nuclei are lumpy, the symmetry planes relevant for collective flow are determined by the geometry of the participating nucleons [39]. They are also called participant planes (PP) or flow planes, and they fluctuate event-by-event, differing for each harmonic order. Because the magnetic field is determined mostly by spectator protons, its direction is expected to be roughly perpendicular to the first-order symmetry plane, which is often called the spectator plane (SP). Thus, the $\psi_{\textsc{RP}}$ in the various terms of Eq. (1) should be replaced by the corresponding-order harmonic planes ($\psi_{n}$). In particular, that for the CME sine term can be replaced by the first-order symmetry plane, which is most closely correlated to the RP. It can also be replaced by the second-order PP angle ($\psi_{2}^{\textsc{pp}}$) because it is closely related to the RP and the most accurately assessed in experiment; in this case, the coefficients $a_{\pm}$ would be reduced accordingly.
@@ 2.2 Charge-dependent physics backgrounds
Because the $\gamma$ correlators, unlike the CME $a_{1}$ parameter in Eq. (1), are parity even and thus subject to QCD physics backgrounds. Charge-independent backgrounds cancel in the $\Delta\gamma$ correlator. However, many physics processes are intrinsically charge dependent; examples are resonance decays and (mini-)jet correlations. There are more contributions to OS pairs than to SS pairs from resonance decays. Charge-ordering is well known in jet fragmentation that also leads to an excess of OS over SS correlations. These short-range charge-dependent correlations can be generally characterized as “clusters”. Local charge conservation (LCC) is preserved in cluster decays, so these background contributions are often referred to as LCC effects.
@@ 2.3 Experimental and analysis methods
Experimentally measured CME observables can contain contributions from both the signal and various background sources. In general, one can write
@@
The background and signal components, although unknown a priori, generally possess different properties. Thus, a central strategy of CME searches is to vary one component while keeping the other approximately fixed, and then study the corresponding change in the observable.
@@ 2.3.2 Event-shape engineering
On the other hand, the statistical fluctuations in the apparent event-by-event anisotropy quantities, e.g. the $v_{2}^{\text{obs}}\equiv\langle{\cos(2\phi-2\psi_{\textsc{EP}})}\rangle$ [74]) and the $q_{2}$ of single particles as well as particle pairs [75, 76], are large. It is tempting to use these quantities calculated from the POIs themselves to select events where the background contribution is small due not only to the small(er) ellipticity of the initial geometry but also to statistical fluctuations in particle distribution in a given event. The so-called event-shape selection (ESS) method is based on this idea. In such selected events the mean $v_{2}$ of POIs is dominated by statistical fluctuations because the event-selection variable is based on POIs. In events where the overall POIs’ $v_{2}=0$, the $v_{2}$ values of the background-contributing sources are not necessarily zero [77]. Such biases are difficult to disentangle, as it would inevitably require the full knowledge of all underlying components of the event. Further quantitative investigations are required to clarify the practical applicability of the ESS method in CME searches.
@@ 2.3.4 Comparative measurements with spectator/participant planes
The elliptic flow values measured with respect to the PP, $v_{2}^{\textsc{pp}}$, and to the SP, $v_{2}^{\textsc{sp}}$, differ by an appreciable 10-20% depending on the collision centrality. This provides an opportunity to extract the CME signal fraction from the two $\Delta\gamma$ measurements with respect to SP and PP, in place of $\psi_{\textsc{rp}}$ in Eq. (2), by [85, 79]
@@ 3.2 Evidence of backgrounds
The first unambiguous experimental evidence of backgrounds dominating the $\Delta\gamma$ observable is from measurements in small systems, where the CME signal is expected to be negligible. Such measurements have been performed in p+Pb collisions by the CMS Collaboration at the LHC [83], and in pp+Au and dd+Au collisions by the STAR Collaboration at RHIC [101]. These results are shown in Fig. 5, where large differences between $\gamma_{\textsc{os}}$} and $\gamma_{\textsc{ss}}$} are observed, comparable to those measured in peripheral heavy ion collisions at similar multiplicities. Because the magnetic field direction is not correlated with the PP direction in small systems, which arises from pure fluctuation effects, any CME signals, even existent in those small-system collisions, are not observable by the $\Delta\gamma$ correlator [83, 84]. The observed $\Delta\gamma$ in small systems and peripheral heavy ion measurements are likely dominated by RP-independent three-particle correlations, as suggested by hijing model calculations [102].
@@ 3.3.1 Event-shape engineering
We note in passing that a large $\eta$ gap between POIs and the region determining $q_{2}$ is useful to mitigate nonflow effects, but it may reduce the power of selecting dynamical fluctuations of $v_{2}$ because of possible longitudinal flow decorrelations. The effects of such decorrelations are, however, expected to be small [109, 110].
It is noteworthy that all ESE analyses assumed that the effects of nonflow and RP-independent three-particle correlation backgrounds are negligible. These effects will have to be accounted for in future high-statistics measurements.
One shortcoming of the ESE method is the extrapolation to zero $v_{2}$, which causes large uncertainties, because of the limited range of dynamical fluctuations of $v_{2}$. It is attempting to exploit statistical fluctuations of $v_{2}$ by using the same POIs as a means to select events. Such an attempt was first reported by the STAR Collaboration in Ref. [74]. Figure 9 left panel presents an example of such an analysis. Unfortunately, such an analysis does not fully eliminate the flow-induced background because the average $v_{2}$’s of background contributing sources are no longer proportional to the statistically fluctuating $v_{2}$ of the POIs used in the analysis [77]. The latest attempt in this ESS approach is to use the $q_{2}$ of pairs of POIs for the event-shape selection. Example results from such an attempt is shown in the right panel of Fig. 9 [76, 71]. The pair $q_{2}$ and $v_{2}$, although directly containing those from the background contributing sources, are dominated by single-particle level contributions from POIs and contains a self-correlation between pairs sharing a common POI [52, 111]. Because the POI pair $q_{2}$ is still selecting on the statistical fluctuations of $v_{2}$, it is challenging, if not at all impossible, to determine how much background remains in the intercept [75, 111]. Experimental results using ESS are therefore not robust without further quantitative studies and modeling.
As a reference, Table 1 contrasts the ESE and ESS methods.
Table 1: Commonalities and contrasts between the ESE and ESS methods, both classifying events by flow-sensitive variables in a narrow centrality bin and projecting the $\Delta\gamma(v_{2})$ as a function of $v_{2}$ to an intercept at $v_{2}=0$ as a CME-sensitive measure.
ESE ESS
POIs same
nonflow in $\Delta\gamma$ same, depending on $\Delta\gamma$ measurement method
event-selection variable $q_{2}$ calculated not from POIs $q_{2}$ calculated from POIs (single particles or pairs)
Nonflow effects from event selection depending on $\eta$ gap between POIs and $q_{2}$ kinematic region likely strong effects because of no $\eta$ gap and self correlations
nature of $v_{2}$ fluctuations dynamical, from fluctuations in initial geometry and/or final-state interactions statistical primarily, inherited from statistical fluctuations in $q_{2}$
statistical precision of intercept poor, because of small dynamical range of $v_{2}$ resulting in data distanced from $v_{2}=0$ high, because of statistical fluctuations covering the $v_{2}=0$ vicinity region
premises and biases background $v_{2}$ strictly proportional to final-state $v_{2}$, so intercept is robust as a CME-sensitive measure; biases may exist due to nonflow effects from event selection background source $v_{2}$ and final-state $v_{2}$, both statistically fluctuating, are not guaranteed to be proportional; interpretation of intercept is unclear
@@ 3.3.2 Isobar collisions
The Ru+Ru to Zr+Zr ratio of the $\Delta\gamma/v_{2}$ variable was observed to be smaller than unity (see Fig. 10), contrary to the expectation of a larger CME signal in Ru+Ru and the same background contributions. The smaller-than-unity ratio results from a few percent difference in the multiplicities, greater (and thus smaller background contribution, see Eq. (9)] in Ru+Ru than Zr+Zr collisions. Such a difference in multiplicities was predicted by energy density functional theory calculations to root in the difference in nuclear structure between the $^{96}_{44}$Ru and $^{96}_{40}$Zr nuclei [86, 114, 115, 116]: the larger number of neutrons in $^{96}_{40}$Zr yields a thicker neutron skin, larger overall size, smaller energy density, and hence lower multiplicity in Zr+Zr collisions. We note that uncertainties in determining nuclear structure from multiplicity or other measurements (mostly in low-energy nuclear reactions) are significant because of large model dependencies. However, the CME background is directly affected by the multiplicity, which is well measured. Inaccuracies in nuclear structure parameters, therefore, do not have a direct bearing on the accuracy of the relative background contributions between the two isobar systems.
@@
Such an estimate of the “baseline” in the isobar comparison has been made by the STAR Collaboration in Refs. [117, 118]. A 2D fitting method of two-particle correlations in $(\Delta\eta,\Delta\phi)$ was used to estimate the (difference) in nonflow contamination in $v_{2}{\{2\}}$ measurements. The three-particle nonflow contamination was estimated by the hijing model following Ref. [119]. Since hijing does not include collective flow, its entire three-particle correlations can be used as an estimate of the RP-independent three-particle correlation background [121, 122]. It was found that the nonflow contribution to $v_{2}$ and the three-particle correlation background were both smaller in Ru+Ru collisions, presumably due to the larger multiplicity dilution in Ru+Ru collisions. The two effects partially cancel each other, and the net effect turns out to be slightly negative compared to the baseline estimated from $r$. These are shown by the brown lines accompanied by shaded bars in Fig. 10 for four of the $\Delta\gamma/v_{2}$ isobar ratio measurements, where the cumulant method is used in measuring $v_{2}{\{2\}}$. The measurements of the $\Delta\gamma/v_{2}$ isobar ratios are consistent with the estimated background baselines, within approximately $1\sigma$ uncertainty. Assuming the magnetic field square difference of 15% between the isobar systems, an upper limit of approximately 10% for $f_{\textsc{cme}}$ is extracted for each of the four measurements at the confidence level of 95%. Nevertheless, we note that the RP-independent three-particle correlations and their systematic uncertainties are estimated by models. For reference, the highest precision data point of the $\Delta\gamma/v_{2}$ isobar ratios in Fig. 10 has a statistical uncertainty of 0.4%, which corresponds to an uncertainty of approximately $3\%$ in terms of $f_{\textsc{cme}}$.
@@ 3.3.3 Spectator/participant planes
As mentioned in Ref. [126], the $f_{\textsc{cme}}^{\text{obs}}$ measurements are still contaminated by nonflow correlations; see Sect. 2.2.2. These nonflow correlations include two-particle nonflow in the $v_{2}{\{2\}}$ measurement and RP-independent three-particle correlations in the $C_{3}$ measurement. Both affect only measurements relative to the PP inferred from TPC particles, not those relative to the SP measured by the ZDCs. The nonflow contribution to $v_{2}{\{2\}}$ reduces $\Delta\gamma\{{\textsc{tpc}}\}$ resulting in an overestimate of the ratio $\Delta\gamma^{\textsc{sp}}/\Delta\gamma^{\textsc{pp}}$, thus a positive contribution to the $f_{\textsc{cme}}^{\text{obs}}$ estimate. The RP-independent three-particle contamination increases $\Delta\gamma\{{\textsc{tpc}}\}$ making its difference from $\Delta\gamma\{{\textsc{zdc}}\}$ smaller, thus gives a negative contribution to the $f_{\textsc{cme}}^{\text{obs}}$ estimate. These two effects partially cancel each other. It is worthwhile to note that, while the experimental measurement of $f_{\textsc{cme}}^{\text{obs}}$ is robust, its interpretation relies on the effect of these nonflow contamination, which is particularly important in this case because of the unconstrained sign of the effect.
The nonflow and the RP-independent three-particle correlation effects in SP/PP measurements have been investigated in Ref. [119] using models. The nonflow effect in $v_{2}$ was estimated by using the ampt model. The effect of RP-independent three-particle correlations is estimated by using the hijing model. The estimated nonflow contribution to $f_{\textsc{cme}}^{\text{obs}}$ is shown in Fig. 12 by the points connected by lines, compared to the STAR measurement. It is found that the net effect of nonflow correlations causes a nearly zero or perhaps negative bias to $f_{\textsc{cme}}$. It will be interesting to have a more data-driven analysis of nonflow effects in the future.
@@
Barring uncertainties in nonflow contamination, the current $f_{\textsc{cme}}^{\text{obs}}$ observed in Au+Au collisions at $\sqrt{s_{{}_{\textsc{nn}}}}=200$ GeV [126] is consistent with the isobar results [81, 117, 118] given the expected an order of magnitude difference in the signal-to-noise ratio between Au+Au and isobar collisions as described in Sect. 3.3.2. It is also consistent with preliminary ESE results in 200 GeV Au+Au collisions at RHIC [106] and with Pb+Pb collisions at the LHC [73, 103], both consistent with zero with relatively large uncertainties as discussed in Sect. 3.3.1. Both the SP/PP and ESE methods are affected by nonflow, in distinctive ways, so more quantitative comparisons need to await nonflow corrections. Nevertheless, there can be non-CME physics mimicking charge separation along the magnetic field, such as a sizable spin alignment of vector mesons or spin correlations of hyperons with respect to the global angular momentum direciton. Such non-CME physics cannot be removed by the SP/PP method, but are an integral part of the extracted signal, as they are practically identical to the CME charge-separation signature.